OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 4 No Solution 0 Any Real Number 1

OpenStudy (anonymous):

Any Real Number

OpenStudy (anonymous):

BECAUSE all real number will satisfy that equation for x

OpenStudy (anonymous):

4x+4=4x+4 4x-4x=4-4 ans is 0